If it takes 5.00 J of work to stretch a Hooke’s-law spring 10.0 cm from its unstressed length, determine the extra work required to stretch it an additional 15.0 cm.


1
Expert's answer
2021-05-30T13:27:43-0400
"W=0.5kx^2\\5=0.5(0.1)^2k\\k=1000\frac{N}{m}\\W'=0.5kx'^2=0.5(1000)(0.15)^2=11.25\ J"